how many molecules of h2o will form with 96.1 g of propane


1
Expert's answer
2021-11-22T02:09:04-0500

Solution:

propane = C3H8

The molar mass of C3H8 is 44.1 g/mol

Hence,

Moles of C3H8 = (96.1 g C3H8) × (1 mol C3H8 / 44.1 g C3H8) = 2.179 mol C3H8 = 2.18 mol C3H8


Balanced chemical equation:

C3H8(g) + 5O2(g) → 3CO2(g) + 4H2O(g)

According to stoichiometry:

1 mol of C3H8 produces 4 mol of H2O

Thus, 2.18 mol of C3H8 produces:

(2.18 mol C3H8) × (4 mol H2O / 1 mol C3H8) = 8.72 mol H2O


One mole of any substance contains 6.02×1023 particles.

Hence,

Number of H2O molecules = (8.72 mol) × (6.02×1023 molecules / 1 mol) = 5.25×1024 molecules

Number of H2O molecules = 5.25×1024 molecules


Answer: 5.25×1024 molecules of H2O will form